Your firm’s discount rate is 8 percent. You are considering the purchase of Truck A or Truck B. Truck A costs $120, has a useful life of 3 years, no salvage value and maintenance costs of $10 per year. Truck B costs $85, has a useful life of 2 years, no salvage value and maintenance costs of $15 per year. Which truck should you select and why?

Truck A because the EAC of -56.6 is less than Truck B’s EAC of -62.7.

Truck B because the EAC of 62.7 is greater than Truck A’s EAC of 56.6.

Truck A because the NPV of 146 if larger than B’s NPV of 112.

Truck A because the EAC of -56.6 is less than Truck B’s EAC of -59.9.
Answer #1

EAC for Turck A = Annual Cost + Purchase cost/Present value annuity factor

= -10 - 120/PVAF(8%, 3 years)

= -10 - 120/2.577

= -$56.6 since outflow

Truck B = -15 - 85/1.7833

= -$62.7

The answer is

Truck A because the EAC of -56.6 is less than Truck B’s EAC of -62.7.